# Ricci almost solitons: complete examples

### v1

@@ 2. The case $D\not=0$
with an additional constant $E$ according to the initial condition $h^{\prime}(0)=E$. Obviously this solution $h^{\prime}(t)$ is globally defined and strictly positive if and only if $D<0$ and $E^{n-1}>-\frac{B}{D}>0$. Recall $B>0$ in any case. In the limit $t\to\infty$ the metric is asymptotic to the product metric $dt^{2}+\big{|}\frac{B}{D}\big{|}^{2/(n-1)}ds_{*}^{2}$ on ×$\mathbb{R}\times M_{*}$ which is Ricci flat. Strictly negative solutions can be considered only if $n$ is even but these cases don’t give any really new solutions for the warped product metric. $\Box$
@@
2. The global shape in the case $D\not=0$ can look like a huge trumpet with a compact Ricci flat space as the fibre. with a cusp at one end and an infinite opening at the other. In the case $D=0$ it looks like a cigar, modelled by the function $h^{\prime}(t)=\sqrt[n-1]{t}$. If $M_{*}$ is compact and Ricci-flat then there is a topological singularity at $t=0$.

### v2

@@ 2. The case $D\not=0$
with an additional constant $E$ according to the initial condition $h^{\prime}(0)=E$. Obviously this solution $h^{\prime}(t)$ is globally defined and strictly positive if and only if $D<0$ and $E^{n-1}>-\frac{B}{D}>0$. Recall $B>0$ in any case. In the limit $t\to\infty$ the metric is asymptotic to the product metric $dt^{2}+\big|\frac{B}{D}\big|^{2/(n-1)}ds_{*}^{2}$ on $\mathbb{R}\times M_{*}$ which is Ricci flat. Strictly negative solutions can be considered only if $n$ is even but these cases don’t give any really new solutions for the warped product metric. $\Box$
@@
2. The global shape in the case $D\not=0$ can look like a huge trumpet with a compact Ricci flat space as the fibre, asymptotically to a cylinder at one end and an infinite opening at the other. In the case $D=0$ it looks like a cigar, modelled by the function $h^{\prime}(t)=\sqrt[n-1]{t}$. If $M_{*}$ is compact and Ricci-flat then there is a topological singularity at $t=0$.
3. In the complete examples with $D\not=0$ in Theorem 4.3 one can compute the functions $f,h^{\prime\prime},h^{\prime\prime\prime}$ explicitely with the positive constants $\alpha:=E^{n-1}+B/D>0,\beta:=-(n-1)D>0,\gamma:=-B/D>0,C>0$ and control their signs. We obtain the following equalities resp. inequalities for all $t\in\mathbb{R}:$
$f(t)$ $=$ $\frac{C}{\beta\gamma}\ln\left(\alpha e^{-\beta t}+\gamma\right)+\frac{C}{\beta}t$
$h^{\prime}(t)$ $=$ $\left(\alpha e^{-\beta t}+\gamma\right)^{\frac{1}{n-1}}>0;h^{\prime\prime}(t)<0;h^{\prime\prime\prime}(t)>0.$
Then the scalar curvature $R=R(t)$ is negative, since by [14, Lem.13]
(17) $h^{\prime 2}(t)R(t)=-\frac{n-2}{n}h^{\prime\prime 2}(t)-\frac{2}{n}h^{\prime}(t)h^{\prime\prime\prime}(t)<0$
for all $t\in\mathbb{R},$ and hence $\lim_{t\to\infty}R(t)=0.$ Also note that by [14, Lem.13]:
$\textrm{Ric}(\partial_{t},\partial_{t})=-(n-1)\frac{h^{\prime\prime\prime}(t)}{h^{\prime}(t)}<0.$
In [4, Thm. 4.3] for different values of $\rho$ the existence and uniqueness of Ricci-Bourguignon solitons (named gradient steady $\rho$-Einstein solitons in [4]) is shown on warped products with $\textrm{Ric}(\partial_{t},\partial_{t})>0$ for some $t.$
4. The same method can be applied to the classification of rotationally symmetric Ricci-Bourguignon solitons only by solving the ODEs in Corollary 4.2 also for $\lambda_{*}>0.$ We do not see a way to achieve that by direct and explicit integration. Presumably one would have to employ other methods. The situation is similar in the case of Bryant’s Ricci solitons in dimensions $n>2$: The solutions exist but not in terms of elementary functions.
Statements and Declarations
The authors have no relevant financial or non-financial interests to disclose.
Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.